205 gram hydrogen sulfide

reacted with 75.0 g oxygen to

produce sulfur and water.

Find:

a) Gram of sulfur formed

b) gram of excess reactant

used up

c) gram of excess reactant

that remained unreacted


2. 295 gram Rubidium Oxide

reacted with 202 gram calcium

cyanide to produce Rubidium

cyanide and Calcium Oxide.

Find

a)gram of Calcium oxide

b) gram of excess reactant

used up

c) gram of excess reactant that

remained unreacted


1
Expert's answer
2021-09-07T02:08:59-0400

Solution (1):

Balanced cheical equation:

2H2S + O2 2S + 2H2O


Determine moles of each reactant:

(205 g H2S) × (1 mol H2S / 34.1 g H2S) = 6.012 mol H2S

(75.0 g O2) × (1 mol O2 / 31.998 g O2) = 2.344 mol O2


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it.

Let's choose O2:

n(H2S) = 2 × n(O2) = (2 × 2.344) = 4.688 mol


The calculation above means that we need 4.688 mol of H2S to completely react with O2.

We have 6.012 mol of H2S and therefore more than enough hydrogen sulfide.

Thus hydrogen sulfide (H2S) is in excess and oxygen (O2) must be the limiting reagent.

Hence, O2 will run out first.


Therefore,

a)

n(S) = 2 × n(O2) = (2 × 2.344) = 4.688 mol

(4.688 mol S) × (32.065 g S / 1 mol S) = 150.32 g S = 150 g S (formed)

b)

n(H2S) = 4.688 mol

(4.688 mol H2S) × (34.1 g H2S / 1 mol H2S) = 159.86 g H2S = 160 g H2S (used up)

c)

205 g H2S - 160 g H2S = 45.0 g H2S (remained unreacted)


Answer (1):

a) 150 grams of sulfur (S) formed.

b) 160 grams of excess reactant (H2S) used up.

c) 45.0 grams of excess reactant (H2S) remained unreacted



Solution(2):

Balanced cheical equation:

Rb2O + Ca(CN)2 → 2RbCN + CaO


Determine moles of each reactant:

(295 g Rb2O) × (1 mol Rb2O / 186.94 g Rb2O) = 1.578 mol Rb2O

(202 g Ca(CN)2) × (1 mol Ca(CN)2 / 92.11 g Ca(CN)2) = 2.193 mol Ca(CN)2


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it.

Let's choose Rb2O:

n(Ca(CN)2) = n(Rb2O) = 1.578 mol


The calculation above means that we need 1.578 mol of Ca(CN)2 to completely react with Rb2O.

We have 2.193 mol of Ca(CN)2 and therefore more than enough calcium cyanide.

Thus calcium cyanide (Ca(CN)2) is in excess and rubidium oxide (Rb2O) must be the limiting reagent.

Hence, Rb2O will run out first.


Therefore,

a)

n(CaO) = n(Rb2O) = 1.578 mol

(1.578 mol CaO) × (56.0774 g CaO / 1 mol CaO) = 88.49 g CaO = 88.5 g CaO (formed)

b)

n(Ca(CN)2) = 1.578 mol

(1.578 mol Ca(CN)2) × (92.11 g Ca(CN)2 / 1 mol Ca(CN)2) = 145.35 g Ca(CN)2 = 145 g Ca(CN)2 (used up)

c)

202 g Ca(CN)2 - 145 g Ca(CN)2 = 57.0 g Ca(CN)2 (remained unreacted)


Answer (2):

a) 88.5 grams of calcium oxide (CaO) formed.

b) 145 grams of excess reactant (Ca(CN)2) used up.

c) 57.0 grams of excess reactant (Ca(CN)2) remained unreacted